# Weak coloring numbers of minor-closed graph classes

### v1

@@ 5. The base case: Graphs with no $\mathcal{F}$-rich model of a given tree
  3. (c’)
    for every component $C$ of $G-S^{\prime}$, $N_{G}(V(C))\subseteq P_{i}\cup P_{j}$ for some $i,j\in\{0,\dots,\ell^{\prime}\}$ and either $i=j$ or $j=i+1$;
@@
    $P_{i}^{\prime}\subseteq R_{i}^{\prime}\subseteq\bigcup_{j\in\{0,\dots,i\}}P_{j}^{\prime}$ for every $i\in\{0,\dots,\ell^{\prime}\}$;
@@
    $\operatorname{wcol}_{r}(G,R_{i}^{\prime})\leqslant(\delta(k,d+1)+3)\cdot r$ for every $i\in[\ell]$ and for every positive integer $r$.
Let $S=U\cup S_{1}\cup(S^{\prime}-U^{\prime})$, $P_{0}=U$, $P_{1}=S_{1}$ and $P_{i}=P^{\prime}_{i-1}$ for every $i\in\{2,\dots,\ell^{\prime}+1\}$ – we set $\ell=\ell^{\prime}+1$. Let $R_{1}=S_{0}\cup V(Q)$ and $R_{i}=R_{i-1}^{\prime}$ for every $i\in\{2,\dots,\ell\}$. We claim that $(P_{0},\dots,P_{\ell})$ is a layering of $S$ in $G$ and (a)-(e) hold, which will complete the proof of the lemma.
Let $i,j\in\{0,\dots,\ell\}$ with $i<j$ and assume that there is an edge a vertex in $P_{i}$ and a vertex in $P_{j}$ in $G$. If $i\geqslant 1$, then $P_{i}\subseteq P^{\prime}_{i-1}$ and $P_{j}=P^{\prime}_{j-1}$, which implies $|i-j|\leqslant 1$ since $(P_{0}^{\prime},\dots,P_{\ell^{\prime}}^{\prime})$ is a layering of $G[S^{\prime}]$. Otherwise, $i=0$ and since $N_{G}(V(G-U^{\prime}))\subseteq S_{1}=P_{1}$, we have $j=1$, which implies that $(P_{0},\dots,P_{\ell})$ is a layering of $S$ in $G$.
Let $F\in\mathcal{F}$. If $V(F)\cap(U\cup S_{1})\neq\emptyset$, then $V(F)\cap S\neq\emptyset$. Otherwise, $F\subseteq G-U$ and $V(F)\cap S_{0}=\emptyset$, and therefore by 17.(a), $F\notin\mathcal{F}_{0}$, so in particular, $V(F)\cap N_{G}(U)=\emptyset$. In this case, $F\in\mathcal{F}^{\prime}$, and by (a’), $V(F)\cap S^{\prime}\neq\emptyset$. This proves (a).
Item (b) holds since $G[U\cup S_{1}]$ is connected and $G[S^{\prime}]$ is connected by (b’).
Let $C$ be a component of $G-S$. Either $C$ is a component of $G-U-S_{1}$ with $C\notin\mathcal{C}_{0}$ or $V(C)\cap U^{\prime}=\emptyset$. In the former case, $N_{G}(V(C))\subseteq U\cup S_{1}=P_{0}\cup P_{1}$, thus, assume the latter case. In particular, $C$ is a component of $G-S^{\prime}$. By (c’), $N_{G}(V(C))$ is included in at most two consecutive parts of $(P^{\prime}_{0},\dots,P^{\prime}_{\ell^{\prime}})$. If $N_{G}(V(C))\subseteq P^{\prime}_{\ell^{\prime}}$, and so $N_{G}(V(C))\subseteq P_{\ell^{\prime}+1}$, or there exists $i\in\{0,\dots,\ell^{\prime}-1\}$ such that $N_{G}(V(C))\subseteq P^{\prime}_{i}\cup P^{\prime}_{i+1}$. If $i=0$, then $N_{G}(V(C))\subseteq U^{\prime}\cup P^{\prime}_{1}$, and since $N_{G}(V(G-U^{\prime}))\subseteq S_{1}$, we have $N_{G}(V(C))\subseteq S_{1}\cup P^{\prime}_{1}=P_{1}\cup P_{2}$. If $i>0$, then $N_{G}(V(C))\subseteq P^{\prime}_{i}\cup P^{\prime}_{i+1}=P_{i-1}\cup P_{i}$. Hence, (c) holds.
Item (d) holds by construction and (d’).
Finally, let $i\in[\ell]$ and let $r$ be a positive integer. If $i=1$, then we set $R_{1}=S_{0}\cup V(Q)$. We have $R_{1}\subseteq P_{0}\cup P_{1}=U\cup S_{1}$ and $P_{1}=S_{1}\subseteq R_{1}$. Moreover,
$\operatorname{wcol}_{r}(G,S_{0}\cup V(Q)))\leqslant\delta(k,d+1)+(2r+1)\leqslant(\delta(k,d+1)+3)\cdot r$
by 17.(c) and Observation 13. If $i>1$, then by (e’), there exists $R^{\prime}_{i-1}\subseteq\bigcup_{a\in\{0,\dots,i-1\}}P^{\prime}_{a}$ with $P^{\prime}_{i-1}\subseteq R^{\prime}_{i-1}$ and $\operatorname{wcol}_{r}(G,R^{\prime}_{i-1})\leqslant(\delta(k,d+1)+3)\cdot r$. We set $R_{i}=R_{i-1}^{\prime}$. This proves that (e) holds and ends the proof of the lemma. ∎
@@
Now we are ready to prove the theorem for graphs with no $\mathcal{F}$-rich model of a fixed star. This part of the argument follows ideas from the proof by [13] that $\operatorname{wcol}_{r}(P)\leqslant 1+\lceil\log r\rceil$ for all paths $P$ – we explain the idea of this proof in Figure 10.
@@
Figure 10. Consider the path $P=p_{0}\dots p_{22}$. Let $r=7$. Then, $s=3=\left\lceil\log(r+1)\right\rceil$. We claim that $\operatorname{wcol}_{r}(P)\leqslant 1+s$. We mark every eigth vertex with $3$ (this corresponds to the set $I_{3}$ in the proof of Lemma 31). Then, we mark each fourth unmarked vertex with $2$ (the set $I_{2}-I_{3}$), every second unmarked vertex with $1$ (the set $I_{1}-I_{2}$), and finally all remaining vertices with 00 (the set $I_{0}-I_{1}$). Next, we construct an ordering on the vertices $p_{0},\dots,p_{22}$. First, preserving the original ordering in the path, we put the vertices marked with $3$, then similarly the ones marked with $2$, with $1$, and with 00. The lower part of the figure depicts this ordering in a slightly convoluted way that we will explain. It suffices to argue that for every $u\in V(P)$, we have $|\operatorname{WReach}_{r}[P,\sigma,u]|\leqslant 4$. We build an auxiliary tree (depicted in green in the figure), where for every $i\in\{3,2,1\}$, we connect every vertex marked with $i$ with the closest (in the path) vertices marked with $i-1$. It is not hard to see that $\operatorname{WReach}_{r}[P,\sigma,u]$ consists only of the ancestors of $u$ in the auxiliary tree.
@@
Let $k,d$ be positive integers. For every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$, then there is a set $S\subseteq V(G)$ such that
@@
Let $G$ be a connected $K_{k}$-minor-free graph, let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$, and suppose that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. Let $r$ be an integer with $r\geqslant 2$. Let $U$ be an arbitrary singleton of a vertex in $G$. Lemma 30 applied to $G$, $\mathcal{F}$, and $U$ gives $S\subseteq V(G)$, a layering $(P_{0},\dots,P_{\ell})$ of $G[S]$ with $P_{0}=U$, and sets $R_{1},\dots,R_{\ell}\subseteq S$ such that
@@
    $P_{i}\subseteq R_{i}\subseteq\bigcup_{j\in\{0,\dots,i\}}P_{j}$ for every $i\in[\ell]$;
@@
    $\operatorname{wcol}_{r}(G,R_{i})\leqslant(\delta(k,d+1)+3)\cdot r$ for every $i\in[\ell]$ and for every positive integer $r$.
Moreover, since $(P_{0},\dots,P_{\ell})$ is a layering of $S$ in $G$ and by 30.(c) we obtain
  1. ($\star$)
    for each $i\in\{0,\dots,\ell\}$, $P_{i}$ separates $\bigcup_{a\in\{0,\dots,i\}}P_{a}$ and $\bigcup_{a\in\{i,\dots,\ell\}}P_{a}$ in $G$.
Note that (a) holds by 30.(a) and (b) holds by 30.(b). In order to conclude the proof, it suffices to show (c).
Let $s=\lceil\log(r+1)\rceil$. For every $i\in\{0,\dots,s\}$, let $I_{i}=\big{\{}j\in\{0,\ldots,\ell\}\mid j=0\mod 2^{i}\big{\}}$. See Figure 10 again. We construct recursively families $\{R^{\prime}_{j}\}_{j\in\{0,\dots,\ell\}}$ and $\{S_{i}\}_{i\in\{0,\dots,s\}}$ of subsets of $V(G)$ and a family $\{\sigma_{j}\}_{j\in\{0,\dots,\ell\}}$ such that $\sigma_{j}$ is an ordering of $R_{j}^{\prime}$. For every $j\in I_{s}$, let
$R_{j}^{\prime}=R_{j}-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}\}}P_{a}$
and let $S_{s}=\bigcup_{j\in I_{s}}R^{\prime}_{j}$. If $j<2^{s}$, then $R_{j}^{\prime}=R_{j}$, and so by 30.(e), $\operatorname{wcol}_{r}(G,R_{j}^{\prime})\leqslant(\delta(k,d+1)+3)\cdot r$. Now assume that $j\geqslant 2^{s}$. By ($\star$)5, $P_{j-2^{s}}$ separates $\bigcup_{a\in\{0,\dots,j-2^{s}\}}P_{a}$ and $\bigcup_{a\in\{j-2^{s},\dots,\ell\}}P_{a}$ in $G$. In particular, $P_{j-2^{s}}$ separates $\bigcup_{a\in\{0,\dots,j-2^{s}\}}P_{a}$ and $R_{j}^{\prime}$ in $G$. Since $P_{j-2^{s}}\subseteq R_{j-2^{s}}^{\prime}$ by 30.(d), by ($\star$)5 and Observation 10, we have
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}\}\cap I_{s}}R^{\prime}_{a},R^{\prime}_{j}\right)$ $=\operatorname{wcol}_{r}\left(G-P_{j-2^{s}},R_{j}^{\prime}\right)$
$=\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}\}}P_{a},R^{\prime}_{j}\right).$
@@
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}\}}P_{a},R^{\prime}_{j}\right)$ $\leqslant\operatorname{wcol}_{r}(G,R_{j})$ by Observation 14
@@
For every $j\in I_{s}$, let $\sigma_{j}$ be an ordering of $R_{j}^{\prime}$ such that
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}\}\cap I_{s}}R^{\prime}_{a},R^{\prime}_{j},\sigma_{j}\right)\leqslant(\delta(k,d+1)+3)\cdot r.$
Next, let $i\in\{0,\dots,s-1\}$ and assume that $S_{i+1}$ is defined. Now for every $j\in I_{i}-I_{i+1}$, let
$R_{j}^{\prime}=\left(R_{j}-\textstyle\bigcup_{a\in\{0,\dots,j-2^{i}\}}P_{a}\right)-S_{i+1},$
and let $S_{i}=\bigcup_{j\in I_{i}}R^{\prime}_{j}$. Let $j\in I_{i}-I_{i+1}$. We have $j-2^{i}\in I_{i+1}$ and therefore $P_{j-2^{i}}\subseteq S_{i+1}$. By ($\star$)5, $P_{j-2^{i}}$ separates $\bigcup_{a\in\{0,\dots,j-2^{i}\}}P_{a}$ and $\bigcup_{a\in\{j-2^{i},\dots,\ell\}}P_{a}$ in $G$. It follows that
$\operatorname{wcol}_{r}(G-S_{i+1},R^{\prime}_{j})=\operatorname{wcol}_{r}(G-P_{j-2^{i}},R^{\prime}_{j})=\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{i}\}}P_{a},R^{\prime}_{j}\right).$
@@
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{i}\}}P_{a},R^{\prime}_{j}\right)$ $\leqslant\operatorname{wcol}_{r}(G,R_{j})$ by Observation 14
@@
Let $\sigma_{j}$ be an ordering of $R_{j}^{\prime}$ such that
@@
This concludes the construction of the families.
Let $S=S_{0}$. The sets $\{R^{\prime}_{j}\}_{j\in\{0,\dots,\ell\}}$ are pairwise disjoint, and they partition $S$. Let $\sigma$ be an ordering of $S$ such that

### v2

@@ 5. The base case: Graphs with no $\mathcal{F}$-rich model of a given tree
    $G[R^{\prime}_{i}]$ is connected for every $i\in\{2,\dots,\ell^{\prime}\}$;
@@
    $W^{\prime}_{i-1}\cap W^{\prime}_{i}\subseteq R^{\prime}_{i}\subseteq\bigcup_{j\in\{0,\dots,i-1\}}W^{\prime}_{j}$ for every $i\in\{2,\dots,\ell^{\prime}\}$;
  6. (f’)
    $W^{\prime}_{i}$ and $W^{\prime}_{i+2}$ are disjoint for every $i\in\{0,\dots,\ell^{\prime}-2\}$; and
  7. (g’)
    $\operatorname{wcol}_{r}(G,R^{\prime}_{i})\leqslant(\delta(k,d+1)+3)\cdot r$ for every $i\in\{2,\dots,\ell\}$ and for every positive integer $r$.
@@
Let $\ell=\ell^{\prime}+1$, $W_{0}=U$, $W_{1}=U^{\prime}$, $W_{2}=(W^{\prime}_{1}-U^{\prime})\cup(S_{1}-U)$, $W_{i}=W^{\prime}_{i-1}$ for every $i\in\{3,\dots,\ell\}$, $R_{2}=S_{1}$, and $R_{i}=R^{\prime}_{i-1}$ for every $i\in\{3,\dots,\ell\}$. Note that (a) holds by construction. We claim that $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$ and (b)-(g) hold, which completes the proof of the lemma.
Let $u\in V(G)$. We claim that $I=\{i\in\{0,\dots,\ell\}\mid u\in W_{i}\}$ is an interval. Since $(W^{\prime}_{0},\dots,W^{\prime}_{\ell^{\prime}})$ is a path decomposition of $G$, $I^{\prime}=\{i\in\{0,\dots,\ell^{\prime}\}\mid u\in W^{\prime}_{i}\}$ is an interval. If $u\not\in U^{\prime}=W^{\prime}_{0}$, then $I=\{i\in\{2,\dots,\ell\}\mid u\in W^{\prime}_{i-1}\}=\{i+1\mid i\in I^{\prime}\}$, which is an interval too. Now suppose that $u\in U^{\prime}$, and so $0\in I^{\prime}$. If $u\not\in S_{1}-U$, then $u\not\in W_{2}$ and $u\not\in W^{\prime}_{i}$ for every $i\in\{2,\dots,\ell^{\prime}\}$ by (a’) and (f’). Hence $I=\{0,1\}$ if $u\in U$, and $I=\{1\}$ otherwise, which is an interval in both cases. If $u\in S_{1}-U$, then $u\not\in U=W_{0}$, and so $I=\{1\}\cup\{i+1\mid i\in I^{\prime}-\{0\}\}$, which is an interval. This proves that $I$ is an interval.
Let $uv$ be an edge of $G$. We claim that there exists $i\in\{0,\dots,\ell\}$ such that $u,v\in W_{i}$. If there exists $i^{\prime}\in\{2,\dots,\ell^{\prime}\}$ such that $u,v\in W^{\prime}_{i^{\prime}}$, then $u,v\in W^{\prime}_{i^{\prime}}=W_{i^{\prime}-1}$ and we are done. Now suppose that $u$ and $v$ are not both in $W^{\prime}_{i}$ for every $i\in\{2,\dots,\ell\}$. Since $(W^{\prime}_{0},\dots,W^{\prime}_{\ell^{\prime}})$ is a path decomposition of $G$, there exists $i^{\prime}\in\{0,1\}$ such that $u,v\in W^{\prime}_{i^{\prime}}$. If $i^{\prime}=0$, then $u,v\in W^{\prime}_{0}=U^{\prime}=W_{1}$. Now suppose that $u$ and $v$ are not both in $W^{\prime}_{0}=U^{\prime}$, and so, in particular, $i^{\prime}=1$. Without loss of generality assume that $v\notin U^{\prime}$. It follows that $v\in W_{1}^{\prime}-U^{\prime}\subseteq W_{2}$. Let $C$ be the component of $v$ in $G-U-S_{1}$. Since $v\not\in U^{\prime}$, $C$ belongs to $\mathcal{C}_{0}$, and so $N_{G}(V(C))\cap U=\emptyset$. It follows that $N_{G}(V(C))\cap U=\emptyset$, and so, $u\in S_{1}-U$. Therefore, $u\in W_{2}$, which concludes the claim. Furthermore, we obtained that $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$.
We now prove (b). Consider $F\in\mathcal{F}$. If $F$ intersects $U$, then $V(F)\cap S\neq\emptyset$ since $U\subseteq S$. If $F$ intersects $S_{1}-U$, then $F$ intersects $W_{1}\cap W_{2}\subseteq S$. Now suppose that $F$ is disjoint from $U\cup S_{1}$. Let $C$ be the component containing $F$ in $G-U-S_{1}$. Since $C$ is disjoint from $S_{0}\subseteq S_{1}$, by 17.(a), $C$ is not a member of $\mathcal{F}_{0}$. This implies that $N_{G}(U)\cap V(C)=\emptyset$, and thus, $C\in\mathcal{C}_{0}$. In particular, $F$ is disjoint from $U^{\prime}$, and so, $F\in\mathcal{F}^{\prime}$. By (b’), $V(F)\cap S^{\prime}\neq\emptyset$, hence, there exists $i\in\{2,\dots,\ell^{\prime}\}$ such that $V(F)$ intersects $W^{\prime}_{i-1}\cap W^{\prime}_{i}$. It follows that $W^{\prime}_{i-1}\cap W^{\prime}_{i}=W_{i}\cap W_{i+1}$ and so $V(F)\cap S\neq\emptyset$. This proves (b).
Let us pause to underline a simple observation that follows directly from the construction, (a’), and (f’):
  1. ($\star$)
    for every $i\in\{3,\dots,\ell\}$, we have $W_{i-1}\cap W_{i}=W_{i-2}^{\prime}\cap W_{i-1}^{\prime}$.
By (c’), $S^{\prime}=U^{\prime}\cup\bigcup_{i\in\{2,\dots,\ell^{\prime}\}}(W^{\prime}_{i-1}\cap W^{\prime}_{i})$ induces a connected subgraph of $G$. In particular, every component of $G[S^{\prime}]-U^{\prime}$ has a neighbor in $N_{G}(V(G)-U^{\prime})\subseteq S_{1}$. Since $G[U\cup S_{1}]$ is connected, it follows that $(S^{\prime}-U^{\prime})\cup U\cup S_{1}$ induces a connected subgraph of $G$. However, $S=(S^{\prime}-U^{\prime})\cup U\cup S_{1}$ by ($\star$)5, which yields (c).
For every $i\in\{3,\dots,\ell\}$, $R_{i}=R^{\prime}_{i-1}$ induces a connected subgraph of $G$ by (d’), and $R_{2}=S_{1}$ induces a connected subgraph of $G$ by definition, hence (d) follows.
For the proof of (e), first, observe that by construction, $\bigcup_{j\in\{0,\dots,i-1\}}W_{j}=\bigcup_{j\in\{0,\dots,i-2\}}W^{\prime}_{j}$ for every $i\in\{2,\dots,\ell\}$. In particular, it follows that $R_{i}=R^{\prime}_{i-1}\subseteq\bigcup_{j\in\{0,\dots,i-2\}}W^{\prime}_{j}=\bigcup_{j\in\{0,\dots,i-1\}}W_{j}$ for every $i\in\{3,\dots,\ell\}$ by (e’). Moreover, $R_{2}=S_{1}\subseteq W_{1}$. It remains to show that $W_{i-1}\cap W_{i}\subseteq R_{i}$ for every $i\in\{2,\dots,\ell^{\prime}\}$. For $i=2$, $W_{1}\cap W_{2}=S_{1}-U\subseteq R_{2}$. For $i\in\{3,\dots,\ell\}$, $W_{i-1}\cap W_{i}=W^{\prime}_{i-2}\cap W^{\prime}_{i-1}\subseteq R^{\prime}_{i-1}=R^{\prime}_{i}$ by ($\star$)5 and (e’). This gives (e).
For every $i\in\{3,\dots,\ell-2\}$, $W_{i}\cap W_{i+2}=W^{\prime}_{i-1}\cap W^{\prime}_{i+1}=\emptyset$ by (f’). Moreover, $W^{\prime}_{3}=W_{4}$ is disjoint from $S_{1}-U\subseteq W^{\prime}_{0}$ by (f’). Hence $W_{2}\cap W_{4}=W^{\prime}_{1}\cap W^{\prime}_{3}=\emptyset$ by (f’). Similarly, $W^{\prime}_{2}=W_{3}$ is disjoint from $U^{\prime}=W^{\prime}_{0}$ by (f’). Hence $W_{1}\cap W_{3}=W^{\prime}_{0}\cap W^{\prime}_{2}=\emptyset$. Finally, $W_{0}\cap W_{2}=\emptyset$ by construction, and so, (f) holds.
It remains to show (g). First, for every $i\in\{3,\dots,\ell\}$, $R_{i}=R^{\prime}_{i-1}$ and so $\operatorname{wcol}_{r}(G,R_{i})\leqslant(\delta(k,d+1)+3)\cdot r$ for every positive integer $r$ by (g’). Moreover, $R_{2}=S_{1}=V(Q)\cup S_{0}$. Hence
$\operatorname{wcol}_{r}(G,R_{2})\leqslant\delta(k,d+1)\cdot r+(2r+1)\leqslant(\delta(k,d+1)+3)\cdot r$
for every positive integer $r$, using 17.(c) and Observation 13. This shows that (g) holds, which concludes the proof of the lemma. ∎
@@
Figure 10. Consider the path $P=p_{1}\cdots p_{22}$. Let $r=7$ and $s=3=\left\lceil\log r\right\rceil$. We claim that $\operatorname{wcol}_{r}(P)\leqslant 2+s$. We mark every eighth vertex with $3$ (this corresponds to the set $I_{3}$ in the proof of Lemma 31). Then, we mark each fourth unmarked vertex with $2$ (the set $I_{2}-I_{3}$), every second unmarked vertex with $1$ (the set $I_{1}-I_{2}$), and finally all remaining vertices with 00 (the set $I_{0}-I_{1}$). Next, we construct an ordering on the vertices $p_{1},\dots,p_{22}$. First, preserving the original ordering in the path, we put the vertices marked with $3$, then similarly the ones marked with $2$, with $1$, and with 00. It suffices to argue that for every $u\in V(P)$, we have $|\operatorname{WReach}_{r}[P,\sigma,u]|\leqslant 5$. We build an auxiliary tree (depicted in green in the figure), where for every $i\in\{3,2,1\}$, we connect every vertex marked with $i$ with the closest vertices in the path marked with $i-1$. It is not hard to see that $\operatorname{WReach}_{r}[P,\sigma,u]$ consists only of the ancestors of $u$ in the auxiliary tree.
Now we are ready to prove Theorem 8 for graphs with no $\mathcal{F}$-rich model of a fixed star. This part of the argument follows ideas from the proof by [13] that $\operatorname{wcol}_{r}(P)\leqslant 2+\lceil\log r\rceil$ for every path $P$ and every positive integer $r$ – we explain the idea of this proof in Figure 10.
@@
Let $k,d$ be positive integers. For every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$, then there is a set $S\subseteq V(G)$ such that
@@
Let $G$ be a connected $K_{k}$-minor-free graph, let $\mathcal{F}$ be a family of connected subgraphs of $G$, and suppose that $G$ has no $\mathcal{F}$-rich model of $F_{2,d}$. Let $r$ be an integer with $r\geqslant 2$. Let $U$ be an arbitrary singleton of a vertex in $G$. Lemma 30 applied to $G$, $\mathcal{F}$, and $U$ gives a path decomposition $(W_{0},\dots,W_{\ell})$ and sets $R_{2},\dots,R_{\ell}\subseteq V(G)$ such that for $S^{\prime}=U\cup\bigcup_{i\in\{2,\dots,\ell\}}(W_{i-1}\cap W_{i})$,
@@
    $G[R_{i}]$ is connected for every $i\in\{1,\dots,\ell\}$;
@@
    $W_{i-1}\cap W_{i}\subseteq R_{i}\subseteq\bigcup_{j\in\{0,\dots,i-1\}}W_{j}$ for every $i\in\{2,\dots,\ell\}$;
  6. 30.(f)
    $W_{i}$ and $W_{i+2}$ are disjoint for every $i\in\{0,\dots,\ell-2\}$; and
  7. 30.(g)
    $\operatorname{wcol}_{r}(G,R_{i})\leqslant(\delta(k,d+1)+3)\cdot r$ for every $i\in\{2,\dots,\ell\}$.
For convenience, we set $R_{1}=U$.
Let $s=\lceil\log(r+1)\rceil$. For every $i\in\{0,\dots,s\}$, let $I_{i}=\{i\in\{1,\dots,\ell\}\mid j=0\mod 2^{i}\}$. We construct recursively families $\{R^{\prime}_{j}\}_{j\in\{1,\dots,\ell\}}$ and $\{S_{i}\}_{i\in\{0,\dots,s\}}$ of subsets of $V(G)$ and a family $\{\sigma_{j}\}_{j\in\{1,\dots,\ell\}}$ such that $\sigma_{j}$ is an ordering of $R_{j}^{\prime}$ for every $j\in\{1,\dots,\ell\}$. For every $j\in I_{s}$, let
$R_{j}^{\prime}=R_{j}-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}-1\}}W_{a}$
and let $S_{s}=\bigcup_{j\in I_{s}}R^{\prime}_{j}$. Let $j\in I_{s}$. If $j<2\cdot 2^{s}$, then $j=2^{s}$ and $R_{j}^{\prime}=R_{j}$, and so by 30.(g), $\operatorname{wcol}_{r}(G,R_{j}^{\prime})\leqslant(\delta(k,d+1)+3)\cdot r$. Now assume that $j\geqslant 2\cdot 2^{s}$. Since $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$, $W_{j-2^{s}-1}\cap W_{j-2^{s}}$ separates $\bigcup_{a\in\{0,\dots,j-2^{s}-1\}}W_{a}$ and $\bigcup_{a\in\{j-2^{s},\dots,\ell\}}W_{a}$ in $G$. Since $W_{j-2^{s}-1}\cap W_{j-2^{s}}\subseteq R_{j-2^{s}}^{\prime}$ (by 30.(e)), by Observation 10, we obtain
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{1,\dots,j-2^{s}\}\cap I_{s}}R^{\prime}_{a},R^{\prime}_{j}\right)$ $=\operatorname{wcol}_{r}\left(G-(W_{j-2^{s}-1}\cap W_{j-2^{s}}),R^{\prime}_{j}\right)$
$=\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}-1\}}W_{a},R^{\prime}_{j}\right).$
@@
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{s}-1\}}W_{a},R^{\prime}_{j}\right)$ $\leqslant\operatorname{wcol}_{r}(G,R_{j})$ by Observation 14
@@
Let $\sigma_{j}$ be an ordering of $R^{\prime}_{j}$ such that
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{1,\dots,j-2^{s}\}\cap I_{s}}R^{\prime}_{a},R^{\prime}_{j},\sigma_{j}\right)\leqslant(\delta(k,d+1)+3)\cdot r.$
Next, let $i\in\{0,\dots,s-1\}$ and assume that $S_{i+1}$ is defined. Now, for every $j\in I_{i}-I_{i+1}$, let
$R^{\prime}_{j}=\left(R_{j}-\textstyle\bigcup_{a\in\{0,\dots,j-2^{i}-1\}}W_{a}\right)-S_{i+1},$
and let $S_{i}=\bigcup_{j\in I_{i}}R^{\prime}_{j}$. Note that $S_{i+1}\subseteq S_{i}$. Also note that for every $j\in I_{i}$, $W_{j-1}\cap W_{j}\subseteq R_{j}^{\prime}$ by 30.(f). Let $j\in I_{i}-I_{i+1}$. We have $j-2^{i}\in I_{i+1}$, and therefore, $W_{j-2^{i}-1}\cap W_{j-2^{i}}\subseteq R_{j-2^{i}}^{\prime}\subseteq S_{i+1}$. Since $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$, $W_{j-2^{i}-1}\cap W_{j-2^{i}}$ separates $\bigcup_{a\in\{0,\dots,j-2^{i}-1\}}W_{a}$ and $\bigcup_{a\in\{j-2^{i},\dots,\ell\}}W_{a}$ in $G$. It follows by Observation 10 that
$\operatorname{wcol}_{r}(G-S_{i+1},R^{\prime}_{j})$ $=\operatorname{wcol}_{r}\left(G-(W_{j-2^{i}-1}\cap W_{j-2^{i}}),R^{\prime}_{j}\right)$
$=\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{i}-1\}}W_{a},R^{\prime}_{j}\right).$
@@
$\operatorname{wcol}_{r}\left(G-\textstyle\bigcup_{a\in\{0,\dots,j-2^{i}-1\}}W_{a},R^{\prime}_{j}\right)$ $\leqslant\operatorname{wcol}_{r}(G,R_{j})$ by Observation 14
@@
Let $\sigma_{j}$ be an ordering of $R^{\prime}_{j}$ such that
@@
We define $S=S_{0}$. Now, it suffices to show that (a)-(c) hold. Since $S^{\prime}\subseteq S$, (a) holds by 30.(b).
Recall that $G[S^{\prime}]$ is connected by 30.(c). Next, let $C$ be a component of $G[R_{j}^{\prime}]$ for some fixed $j\in I_{s}$. If $V(C)\cap(W_{j-1}\cap W_{j})\neq\emptyset$, then $V(C)\cap S^{\prime}\neq\emptyset$, and so, $G[S^{\prime}\cup V(C)]$ is connected. Thus, assume that $V(C)\cap(W_{j-1}\cap W_{j})=\emptyset$. However, $W_{j-1}\cap W_{j}\subseteq R_{j}^{\prime}$, hence, $C$ has a neighbor in $W_{j-2^{s}-1}$, in particular, in $W_{j-2^{s}-1}\cap W_{j-2^{s}}\subseteq S^{\prime}$. Hence, again $G[S^{\prime}\cup V(C)]$ is connected. In particular, we have just proved that $G[S^{\prime}\cup S_{s}]$ is connected. Next, suppose that $G[S^{\prime}\cup S_{i+1}]$ is connected for some $i\in\{0,\dots,s-1\}$. Let $C$ be a component of $G[R_{j}^{\prime}]$ for some fixed $j\in I_{i}$. If $V(C)\cap(W_{j-1}\cap W_{j})\neq\emptyset$, then $V(C)\cap S^{\prime}\neq\emptyset$, and so, $G[S^{\prime}\cup V(C)]$ is connected. Thus, assume that $V(C)\cap(W_{j-1}\cap W_{j})=\emptyset$. However, $W_{j-1}\cap W_{j}\subseteq R_{j}^{\prime}$, hence, $C$ has a neighbor in $W_{j-2^{i}-1}\cup S_{i+1}$, in particular, in $(W_{j-2^{i}-1}\cap W_{j-2^{i}})\cup S_{i+1}\subseteq S^{\prime}\cup S_{i+1}$. Hence, $G[S^{\prime}\cup S_{i+1}\cup V(C)]$ is connected. Finally, $G[S^{\prime}\cup S_{0}]=G[S]$ is connected, which yields (b).
The sets $\{R^{\prime}_{j}\}_{j\in\{2,\dots,\ell\}}$ are pairwise disjoint, and they partition $S$. Let $\sigma$ be an ordering of $S$ such that